### v3

Abstract
Recent studies have demonstrated the great potential of reduced order modeling for parametric dynamical systems using low-rank tensor decompositions (LRTD). In particular, within the framework of interpolatory tensorial reduced order models (ROM), LRTD is computed for tensors composed of snapshots of the system’s solutions, where each parameter corresponds to a distinct tensor mode. This approach requires full sampling of the parameter domain on a tensor product grid, which suffers from the curse of dimensionality, making it practical only for systems with a small number of parameters. To overcome this limitation, we propose a sparse sampling of the parameter domain, followed by a low-rank tensor completion. The resulting specialized tensor completion problem is formulated for a tensor of order $C+D$, where $C$ fully sampled modes correspond to the snapshot degrees of freedom, and $D$ partially sampled modes correspond to the system’s parameters. To address this non-standard tensor completion problem, we introduce a low-rank tensor format called the hybrid tensor train. Completion in this format is then integrated into an interpolatory tensorial ROM. We demonstrate the effectiveness of both the completion method and the ROM on several examples of dynamical systems derived from finite element discretizations of parabolic partial differential equations with parameter-dependent coefficients or boundary conditions.
1 Introduction
In this paper, we focus on reduced order modeling for a multiparameter dynamical system with a sparse sampling of the parameter domain. A particular setup that we study here is given by a dynamical system
$\mathbf{u}_{t}=F(t,\mathbf{u},\boldsymbol{\alpha}),\quad t\in(0,T),\quad\text{and}~\mathbf{u}|_{t=0}=\mathbf{u}_{0},$ (1)
where a known continuous flow field $F:(0,T)\times\mathbb{R}^{M_{1}}\times\mathcal{A}\to\mathbb{R}^{M_{1}}$ depends on the vector of parameters $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{D})^{T}$ from a parameter domain $\mathcal{A}\subset\mathbb{R}^{D}$. A possible example of (1) is a system of ODEs for nodal values of the finite volume or finite element solution to a parabolic PDE problem, where material coefficients, body forces, boundary conditions, the computational domain (via a mapping into a reference domain), etc., are parameterized by $\boldsymbol{\alpha}$.
After discretization in time, each trajectory $\mathbf{u}=\mathbf{u}(t,\boldsymbol{\alpha}):[0,T)\to\mathbb{R}^{M_{1}}$ is represented by the collection of snapshots $\mathbf{u}(t_{k},\boldsymbol{\alpha})\in\mathbb{R}^{M_{1}}$ at times $0\leq t_{1}<\dots<t_{M_{2}}\leq T$. The snapshots can be organized in a matrix
$\Phi(\boldsymbol{\alpha})=\big[\mathbf{u}(t_{1},\boldsymbol{\alpha}),\dots,\mathbf{u}(t_{M_{2}},\boldsymbol{\alpha})\big]\in\mathbb{R}^{M_{1}\times M_{2}}.$ (2)
Assuming that the parameter domain $\mathcal{A}$ is the $D$-dimensional box
$\mathcal{A}={\textstyle\bigotimes\limits_{i=1}^{D}}[\alpha_{i}^{\min},\alpha_{i}^{\max}],$ (3)
we introduce a Cartesian grid on $\mathcal{A}$ by distributing $K_{i}$ nodes $\{\widehat{\alpha}_{i}^{j}\}_{j=1,\dots,K_{i}}$ within each of the intervals $\alpha_{i}^{\min},\alpha_{i}^{\max}]$ for $i=1,\dots,D$. The nodes of the grid form the set
${\widehat{\mathcal{A}}}=\left\{\widehat{\boldsymbol{\alpha}}=(\widehat{\alpha}_{1},\dots,\widehat{\alpha}_{D})^{T}:\,\widehat{\alpha}_{i}\in\{\widehat{\alpha}_{i}^{k}\}_{k=1}^{K_{i}},~i=1,\dots,D\right\},\quad K=\prod_{i=1}^{D}K_{i}.$ (4)
[…]
$(\boldsymbol{\Phi})_{:,:,k_{1},\dots,k_{D}}=\Phi(\widehat{\boldsymbol{\alpha}}),\quad\widehat{\boldsymbol{\alpha}}=\big(\widehat{\alpha}_{1}^{k_{1}},\dots,\widehat{\alpha}_{D}^{k_{D}}\big)^{T},$ (5)
which is a tensor of order $2+D$ and size $M_{1}\times M_{2}\times K_{1}\times\dots\times K_{D}$. Observe that the first and second indices of the snapshot tensor $\boldsymbol{\Phi}$ are reserved for modes corresponding to the spatial and temporal degrees of freedom, respectively. Depending on any additional tensor structure exhibited by the state variable $\mathbf{u}$, more than two modes of $\boldsymbol{\Phi}$ may be allocated for space-time indexing. In general, we assume $C\geq 1$ space-time modes in $\boldsymbol{\Phi}\in\mathbb{R}^{M_{1}\times\dots\times M_{C}\times K_{1}\times\dots\times K_{D}}$.
It is clear that for a high dimension $D$ or a fine parameter grid $^{\widehat{\mathcal{A}}}$, working directly with $\boldsymbol{\Phi}$ is prohibitively expensive, if possible at all. Therefore, one approach to model order reduction for a system like (1) is based on the assumption that $\boldsymbol{\Phi}$ can be effectively approximated by a low-rank tensor $\widetilde{\boldsymbol{\Phi}}$ in one of the commonly used formats. Depending on the definition of tensor rank, various tensor decomposition formats can be employed, including CANDECOMP/PARAFAC (CP), Tucker (HOSVD), or Tensor Train (TT). While known rigorous analyses revealing how tensor ranks depend on the properties of a parameterized differential equation and the targeted accuracy of the solution manifold recovery, or its smoothness, are very limited (see, e.g., bachmayr2017kolmogorov and discussions in khoromskij2011tensor ; nouy2017low ; bachmayr2023low ), numerical evidence suggests that for many practical parametric PDEs, solutions to such problems or statistics derived from them are well approximated in low-rank formats kressner2011low ; ballani2015hierarchical ; dolgov2015polynomial ; ballani2016reduced ; eigel2017adaptive ; bachmayr2018parametric ; dolgov2019hybrid ; glau2020low ; mamonov2022interpolatory ; mamonov2024tensorial .
Since entry-wise assembly of the snapshot tensor $\boldsymbol{\Phi}$ may not be feasible for large $D$, a more practical approach for finding $\widetilde{\boldsymbol{\Phi}}$ is by solving a tensor completion problem. A general low-rank tensor completion problem can be formulated as finding a minimal rank tensor $\widetilde{\boldsymbol{\Phi}}\in\mathbb{R}^{M_{1}\times\dots\times M_{C}\times K_{1}\times\dots\times K_{D}}$ that fits the tensor $\boldsymbol{\Phi}\in\mathbb{R}^{M_{1}\times\dots\times M_{C}\times K_{1}\times\dots\times K_{D}}$ for a subset of its observed entries:
$\widetilde{\boldsymbol{\Phi}}=\underset{\boldsymbol{\Psi}\in\mathbb{R}^{M_{1}\times\dots\times M_{C}\times K_{1}\times\dots\times K_{D}}}{\operatorname{argmin}}\mbox{rank}(\boldsymbol{\Psi}),\quad\text{s.t.}\quad\boldsymbol{\Psi}|_{\Omega}=\boldsymbol{\Phi}|_{\Omega},$ (6)
where $\boldsymbol{\Psi}|_{\Omega}$ is a restriction of the tensor on the set of indices $\Omega$ of observed entries. The exact fitting in (6) can be relaxed to the approximate one, yielding the inexact completion problem
$\widetilde{\boldsymbol{\Phi}}=\underset{\boldsymbol{\Psi}\in\mathbb{R}^{M_{1}\times\dots\times M_{C}\times K_{1}\times\dots\times K_{D}}}{\operatorname{argmin}}\mbox{rank}(\boldsymbol{\Psi}),\quad\text{s.t.}~\|\boldsymbol{\Psi}|_{\Omega}-\boldsymbol{\Phi}|_{\Omega}\|\leq\varepsilon,$ (7)
with a prescribed $\varepsilon\geq 0$. The inexact completion (7) is a commonly used problem setup for the case of noisy data candes2010matrix and this is the formulation we are interested here. For CP, Tucker, and TT tensor ranks, the completion problems (6) and (7) are NP-hard. Building on the success of solving low-rank matrix completion problems candes2010power ; cai2010singular ; candes2012exact ; vandereycken2013low ; jain2013low ; nguyen2019low , a popular approach is to relax (6)–(7) into convex optimization problems signoretto2010nuclear ; gandy2011tensor ; bengua2017efficient . Other approaches to approximate tensor completion include ALS methods yokota2016smooth ; grasedyck2019stable , Riemannian optimization kressner2014low ; steinlechner2016riemannian , Bayesian methods zhao2015bayesian ; zhao2015bayesian2 , and projection methods rauhut2015tensor , with applications ranging from video recovery to seismic data reconstruction bengua2017efficient ; liu2022efficient .
While the completion problem formulations (6)–(7) accommodate a wide range of applications, model order reduction for systems like (1) imposes additional structure on the set $\Omega$. In particular, we focus on the case where $\Omega$ corresponds to the so-called slice sampling. As indicated by the notation above, in slice sampling, the set of all $C+D$ tensor modes is divided into two groups. The first $C$ modes, $M_{1}\times\dots\times M_{C}$, are fully sampled, while the remaining $D$ modes, $K_{1}\times\dots\times K_{D}$, are sampled sparsely. For a fixed set of $D$ indices $k_{1},\dots,k_{D}$ corresponding to sparsely sampled modes, we define the tensor
$(\boldsymbol{\Phi})_{:,\ldots,:,k_{1},\dots,k_{D}}\in\mathbb{R}^{M_{1}\times\dots\times M_{C}}$ (8)
as a slice, hence the name slice sampling. For the problem (1), we have $C=2$, and in general we assume that $C$ is small so that using Cartesian grid in physical variables is feasible. The slice (8) coincides with the snapshot matrix (2), where $\boldsymbol{\alpha}=\big(\widehat{\alpha}_{1}^{k_{1}},\dots,\widehat{\alpha}_{D}^{k_{D}}\big)^{T}$. The sampled indices $k_{1},\dots,k_{D}$ form a sparse subset of the grid $^{\widehat{\mathcal{A}}}$.
To efficiently exploit slice sampling and improve the computational cost of solving the completion problem (7), we introduce a custom low-rank tensor format called the hybrid tensor train (HTT). Completion in the HTT format involves projecting the slice-sampled tensor onto reduced orthogonal bases along the first $C$ modes (similar to HOSVD), followed by multiple completions of smaller tensors in TT format. These smaller completions can be performed in parallel using existing tensor completion algorithms. The HTT format is particularly well-suited for constructing a Galerkin reduced-order model (ROM) known in the literature as the tensorial ROM (TROM) mamonov2022interpolatory ; mamonov2023analysis ; mamonov2024tensorial , a natural extension of the POD ROM for parametric dynamical systems.
The application of tensor completion methods to reduced order modeling of parametric ODEs or PDEs is rare, and this study aims to explore this direction. In general, the use of tensor methods for solving parametric PDEs is not new. Several studies have developed sparse tensorized solvers for certain high-dimensional and stochastic PDEs  schwab2011sparse ; khoromskij2011tensor ; dolgov2015polynomial ; garreis2017constrained ; nouy2017low ; dolgov2018direct ; dolgov2019hybrid .
The reconstruction of scalar output quantities of parametric solutions in tensor format from incomplete observations was addressed in ballani2015hierarchical ; glau2020low . In ballani2015hierarchical the authors employed a tensor cross approximation, while glau2020low applied TT-completion via Riemannian optimization to recover an option pricing statistic from solutions of parametrized Heston and multi-dimensional Black-Scholes models. Additionally, a comparison of TT-cross interpolation and TT-completion for a parameterized diffusion equation in steinlechner2016riemannian demonstrated that TT-completion requires fewer PDE solver executions to find a low-rank approximation of a particular solution functional.
While the works ballani2015hierarchical ; steinlechner2016riemannian ; glau2020low focused on recovering scalar solution statistics in tensor format, here we aim to approximate a tensor of solution snapshots for subsequent use in building TROM. Similarly to the previous studies  mamonov2022interpolatory ; mamonov2023analysis ; mamonov2024tensorial , we seek a low-rank approximation of a tensor. However, here we compute such an approximation from a sparse sampling of the tensor. Due to the relatively high separation ranks and large sizes of the space and time modes in the snapshot tensor, applying existing completion algorithms in standard low-rank formats is computationally prohibitive. This motivates the introduction of tensor completion in the HTT format customized specifically for slice sampled tensors arising in TROM construction.
The remainder of the paper is organized as follows: Section 2 provides a more detailed explanation of slice sampling and introduces the proposed completion method. We deviate from the traditional approach of defining tensor rank before formulating the completion problem, as we find it more instructive to first explain the method of obtaining the fitting tensor. The resulting rank-revealing format becomes more intuitive afterward. Section 3 describes the Galerkin ROM for the dynamical system. This ROM utilizes HTT as the dimension reduction technique (in place of the standard POD), and we refer to it as HTT-ROM. Section 4 presents the results of numerical experiments.
2 Slice sampling tensor completion
We consider here a problem that we refer to hereafter as slice sampling low-rank tensor completion. Consider a tensor $\boldsymbol{\Phi}$ of order $C+D$ of size $M_{1}\times\ldots\times M_{C}\times K_{1}\times\ldots\times K_{D}$ and two sets of indices
$\Omega_{C}=\bigotimes\limits_{i=1}^{C}\{1,\ldots,M_{i}\},\quad\Omega_{D}=\bigotimes\limits_{j=1}^{D}\{1,\ldots,K_{j}\},$ (9)
where the products are understood in Cartesian sense. The set $\Omega_{C}$ is the set of sliced indices and $\Omega_{D}$ is the set of sampled indices. The sampling (or training) set is the subset
$\widetilde{\Omega}_{D}\subset\Omega_{D}.$ (10)
In what follows we employ multi-indices
$\mathbf{i}=(i_{1},\ldots,i_{C}),\quad\mathbf{j}=(j_{1},\ldots,j_{D}),$ (11)
for the first $C$ and last $D$ indices of $\boldsymbol{\Phi}$, respectively. Then, given the data
[…]
we seek a completing tensor $\widetilde{\boldsymbol{\Phi}}$ solving (6) or (7) with $\Omega=\Omega_{C}\otimes\widetilde{\Omega}_{D}$.
Examining the definition of the index sets (9)–(10), we observe that the term slice sampling refers to the setup in which the first $C$ modes of $\boldsymbol{\Phi}$, the space–time modes of the snapshot tensor, are fully sampled, while the last $D$ modes are only partially sampled. In other words, we learn the tensor $\boldsymbol{\Phi}$ through $C$-dimensional slices. This distinguishes the slice sampling completion problem considered here from conventional low-rank tensor completion settings, in which none of the tensor modes are fully sampled.
Given that the two index sets (9) of entries of $\boldsymbol{\Phi}$ are sampled differently, slice sampling tensor completion is a two-stage process. First, orthonormal bases for the low-dimensional subspaces of the fully sampled modes are computed. Second, TT-completion is performed component-wise in the subspaces computed in the first stage. Finally, the completed tensor $\widetilde{\boldsymbol{\Phi}}$ can be assembled, or one may use the reduced bases and component-wise TT-completions from the first two stages to operate on $\widetilde{\boldsymbol{\Phi}}$ without assembling it explicitly.
2.1 Reduced basis for fully sampled space–time modes
As mentioned above, the first stage of slice sampling completion is the computation of the orthonormal bases for the reduced subspaces corresponding to space–time modes of $\boldsymbol{\Phi}$. Since the first $C$ modes are fully sampled, one may employ a simple technique based on truncated SVD to compute the bases of interest.
First, for all multi-indices in the set $\widetilde{\Omega}_{D}$ introduce a linear enumeration so that it can be written as
$\widetilde{\Omega}_{D}=\left\{\left.\mathbf{j}^{(k)}\in\Omega_{D}\;\right|\;k=1,\ldots,P\right\},\quad P=|\widetilde{\Omega}_{D}|.$ (13)
Consider a tensor $\boldsymbol{\Phi}^{P}$ of order $C+1$ of size $M_{1}\times\ldots\times M_{C}\times P$ with entries
$\Phi^{P}_{\mathbf{i},k}=\Phi_{\mathbf{i},\mathbf{j}^{(k)}},\quad k=1,\ldots,P.$ (14)
Then, assemble $C$ matrices $\mathrm{F}^{(i)}\in\mathbb{R}^{M_{i}\times{M}_{i}^{\prime}P}$, $i=1,\ldots,C$, with $M_{i}^{\prime}=\prod\limits_{{q=1}\atop{q\neq i}}^{C}M_{q}$, defined as
$\mathrm{F}^{(i)}=\mbox{unfold}_{i}\Big(\boldsymbol{\Phi}^{P}\Big),$ (15)
where $\mbox{unfold}_{i}$ denotes the $i^{\text{th}}$-mode unfolding of a tensor. Next, compute the singular value decompositions
$\mathrm{F}^{(i)}=\mathrm{U}^{(i)}\Sigma^{(i)}(\mathrm{V}^{(i)})^{T},\quad i=1,\ldots,C,$ (16)
where each $\Sigma^{(i)}$ contains the singular values $\sigma^{(i)}_{1}\geq\ldots\geq\sigma^{(i)}_{M_{i}}$. Choose a threshold $\varepsilon_{C}\geq 0$ and determine the ranks
$q_{i}=\min\Big\{q~\Big|~\sum_{j>q}\big(\sigma^{(i)}_{j}\big)^{2}\leq\varepsilon_{C}\big\|\mathrm{F}^{(i)}\big\|_{F}^{2}\Big\},\quad i=1,\ldots,C.$ (17)
Using Matlab notation, we denote the matrix containing the first $q_{i}$ columns of $\mathrm{U}^{(i)}$ by
$\widetilde{\mathrm{U}}^{(i)}=\mathrm{U}^{(i)}_{:,1:q_{i}}\in\mathbb{R}^{M_{i}\times q_{i}},\quad i=1,\ldots,C.$ (18)
The columns of matrices $\widetilde{\mathrm{U}}^{(i)}$ comprise the orthonormal bases that we refer to as the reduced bases for all fully sampled modes, $i=1,\ldots,C$.
2.2 Tensor-train completion for partially sampled modes
Once the reduced bases for the fully sampled space–time modes are computed, there exist two options for completion along the partially sampled ones. In order to describe both we recall the definition of $n$-mode tensor-matrix product. Given a tensor $\boldsymbol{\Psi}$ of order $d$ of size $m_{1}\times\ldots\times m_{d}$ and a matrix $\mathrm{M}\in\mathbb{R}^{m\times m_{n}}$, the $n$-mode product $\boldsymbol{\Psi}\times_{n}\mathrm{M}$ is a tensor of order $d$ of size $m_{1}\times m_{n-1}\times m\times m_{n+1}\times m_{d}$ with entries given by
$\boldsymbol{\Psi}\times_{n}\mathrm{M}]_{i_{1},\ldots,i_{n-1},j,i_{n+1},\ldots,i_{d}}=\sum_{k=1}^{m_{n}}\boldsymbol{\Psi}_{i_{1},\ldots,i_{n-1},k,i_{n+1},\ldots,i_{d}}\mathrm{M}_{jk},\quad j=1,\ldots,m.$ (19)
Consider the projected tensor
$\boldsymbol{\Phi}^{q}=\boldsymbol{\Phi}\times_{1}\left(\widetilde{\mathrm{U}}^{(1)}\right)^{T}\times_{2}\left(\widetilde{\mathrm{U}}^{(2)}\right)^{T}\times_{3}\ldots\times_{C}\left(\widetilde{\mathrm{U}}^{(C)}\right)^{T}$ (20)
of order $C+D$ and of size $q_{1}\times\ldots\times q_{C}\times K_{1}\times\ldots\times K_{D}$. With the reduced bases of fully sampled modes at hand, one may compute from the original data $\mathcal{D}$ the projected data
[…]
where
$\Omega_{C}^{q}=\bigotimes\limits_{i=1}^{C}\{1,\ldots,q_{i}\}.$ (22)
Then, the key step in finding $\widetilde{\boldsymbol{\Phi}}$ from $\mathcal{D}$ is to determine $\widetilde{\boldsymbol{\Phi}}^{q}$, a low-rank completion of $\boldsymbol{\Phi}^{q}$ from the projected data $\mathcal{D}^{q}$. Once $\widetilde{\boldsymbol{\Phi}}^{q}$ is found, the completion $\widetilde{\boldsymbol{\Phi}}$ of $\boldsymbol{\Phi}$ is given by
$\widetilde{\boldsymbol{\Phi}}=\widetilde{\boldsymbol{\Phi}}^{q}\times_{1}\widetilde{\mathrm{U}}^{(1)}\times_{2}\widetilde{\mathrm{U}}^{(2)}\times_{3}\ldots\times_{C}\widetilde{\mathrm{U}}^{(C)}.$ (23)
One can easily check [^1] To verify the result in (24), we note that $\|\boldsymbol{\Phi}|_{\Omega}-\widetilde{\boldsymbol{\Phi}}|_{\Omega}\|_{F}=\|\boldsymbol{\Phi}^{P}-\widetilde{\boldsymbol{\Phi}}^{P}\|_{F}$, $\|\boldsymbol{\Phi}|_{\Omega}\|_{F}=\|\boldsymbol{\Phi}^{P}\|_{F}$ and apply (de2000multilinear, , Property 10) after observing that $\widetilde{\boldsymbol{\Phi}}^{P}$ is a truncated HOSVD of $\boldsymbol{\Phi}^{P}$. that $\widetilde{\boldsymbol{\Phi}}$ satisfies
$\big\|\boldsymbol{\Phi}|_{\Omega}-\widetilde{\boldsymbol{\Phi}}|_{\Omega}\big\|_{F}\leq\sqrt{C}\varepsilon_{C}\big\|\boldsymbol{\Phi}|_{\Omega}\big\|_{F},$ (24)
with $\varepsilon_{C}$ from (17). For $\varepsilon_{C}>0$ this renders our completion inexact as in (7).
We still need to address the completion problem for a sliced-sampled tensor, but with reduced dimensions of the slices. As mentioned previously, there are two ways of solving this problem that we consider below.
The first option is to find the whole tensor $\widetilde{\boldsymbol{\Phi}}^{q}$ in a common low-rank tensor format such as TT. Further in the paper we work with with low-rank completion in TT format, but it is certainly possible to use other LRTDs, e.g., in CP or Tucker formats, as well. Finding $\widetilde{\boldsymbol{\Phi}}^{q}$ can be achieved using an existing method for low-rank tensor completion, e.g., stable ALS method from grasedyck2019stable . The main disadvantage of such approach applied to, e.g., model order reduction for parametric dynamical systems, is its high cost both in terms of memory and computational time. To resolve it, we suggest a different approach introduced as the second option below.
The second option is to perform a low-rank TT completion component-wise in the following sense. Let $Q=\left|\Omega_{C}^{q}\right|=\prod_{i=1}^{C}q_{i}$. Consider $Q$ data sets
$\mathcal{D}^{q}_{\mathbf{i}}=\left\{\Phi^{q}_{\mathbf{i},\mathbf{j}}\;\left|\;\forall\mathbf{j}\in\widetilde{\Omega}_{D}\right.\right\},\quad\mathbf{i}\in\Omega_{C}^{q}.$ (25)
Then, we perform $Q$ low-rank TT completions with data $\mathcal{D}^{q}_{\mathbf{i}}$ for each $\mathbf{i}\in\Omega_{C}^{q}$ to obtain TT-tensors
$\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}=\sum_{j_{0}=1}^{r_{0}^{\mathbf{i}}}\cdots\sum_{j_{D}=1}^{r_{D}^{\mathbf{i}}}\mathbf{g}^{\mathbf{i}}_{1,j_{0},j_{1}}\circ\cdots\circ\mathbf{g}^{\mathbf{i}}_{D,j_{D-1},j_{D}},\quad\mathbf{i}\in\Omega_{C}^{q},$ (26)
where $\mathbf{g}^{\mathbf{i}}_{k,j_{k-1},j_{k}}\in\mathbb{R}^{N_{k}}$ for $k=1,\ldots,D$, and for all values of $j_{0},\ldots,j_{D}$ that appear in (26). The summation indices $r_{k}^{\mathbf{i}}$ are the TT compression ranks, where we follow the convention $r_{0}^{\mathbf{i}}=r_{D}^{\mathbf{i}}=1$ for notation convenience.
After computing the completions (26), the entries of the tensor $\widetilde{\boldsymbol{\Phi}}^{q}$ are simply
$\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i},\mathbf{j}}=\left[\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}\right]_{\mathbf{j}},\quad\mathbf{i}\in\Omega_{C}^{q},\quad\mathbf{j}\in\Omega_{D}.$ (27)
Then, the completion $\widetilde{\boldsymbol{\Phi}}$ of $\boldsymbol{\Phi}$ can be computed via (23).
Summarizing, the HTT format we introduced to complete a sliced-sampled tensor is given by $C$ matrices and $Q$ tensors in TT format:
$\rm{HTT}\big(\widetilde{\boldsymbol{\Phi}}\big)=\left\{\widetilde{\mathrm{U}}^{(i)}\in\mathbb{R}^{M_{i}\times q_{i}},\quad i=1,\ldots,C,\quad\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}\in\mathbb{R}^{K_{1}\times\dots\times K_{D}},~\mathbf{i}\in\Omega_{C}^{q}\right\}.$ (28)
We call $\{q_{i}\}_{i=1,\dots,C}$ the C-ranks and $\{r^{\mathbf{i}}_{i}\}_{i=0,\dots,D}$ the D-ranks of $\widetilde{\boldsymbol{\Phi}}$.
Assuming for simplicity that $M_{i}=M$, $K_{j}=K$, all C-ranks are equal to $q$ and all D-ranks are equal to $r$, the representation complexity of $\boldsymbol{\Phi}$ is $CMq+(q^{C})D(r^{2})K$. We observe that the complexity grows exponentially in $C$. However, in applications we are interested in, $C$ is typically small. In Section 3, we provide the details of the case when the slice sampling completion is applied to build a ROM for a general parametric dynamical system with $C=2$.
2.3 Algorithm for slice sampling tensor completion
We summarize the discussion of slice sampling tensor completion in HTT format in the algorithm below.
Algorithm 1 (Slice sampling tensor completion in HTT format)
Input: Index set $\widetilde{\Omega}_{D}$ as in (10) and the corresponding data $\mathcal{D}$ as in (12), threshold $\varepsilon_{C}\geq 0$.
  1. 1.
    Form the tensor $\boldsymbol{\Phi}^{P}$ from the data $\mathcal{D}$ as in (14).
  2. 2.
    For $i=1,\ldots,C$:
    • (a)
      Form the unfolding $\mathrm{F}^{(i)}=\mbox{unfold}_{i}\big(\boldsymbol{\Phi}^{P}\big)$;
    • (b)
      Compute the SVD of the unfolding matrix $\mathrm{F}^{(i)}=\mathrm{U}^{(i)}\Sigma^{(i)}(\mathrm{V}^{(i)})^{T}$;
    • (c)
      Choose the rank $q_{i}$ using (17) and form the matrix $\widetilde{\mathrm{U}}^{(i)}=\mathrm{U}^{(i)}_{:,1:q_{i}}$, the columns of which form the reduced basis.
  3. 3.
    For each multi-index $\mathbf{i}\in\Omega^{q}_{C}$:
    • (a)
      Compute the projected data
      $\mathcal{D}^{q}_{\mathbf{i}}=\left\{\Phi_{\mathbf{i},\mathbf{j}^{(k)}}\times_{1}\big(\widetilde{\mathrm{U}}^{(1)}\big)^{T}\times_{2}\big(\widetilde{\mathrm{U}}^{(2)}\big)^{T}\times_{3}\ldots\times_{C}\big(\widetilde{\mathrm{U}}^{(C)}\big)^{T}\right\}_{k=1}^{P},$ (29)
      where the multi-indices $\mathbf{j}^{(k)}$ are as in (13);
    • (b)
      Perform low-rank completion in TT format (26) with data $\mathcal{D}^{q}_{\mathbf{i}}$ to obtain TT tensor $\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}$ with compression ranks $r^{\mathbf{i}}_{k}$, $k=0,1,\ldots,D$.
[…]
2.4 An estimate for the $\{C,D\}$-ranks
Let us prove one estimate of the $C$- and $D$-ranks of a tensor through the ranks of its unfoldings. To this end, denote by $\mbox{matr}_{k}(\mathbf{A})\in\mathbb{R}^{(K_{1}\dots K_{k})\times(K_{k+1}\dots K_{N})}$ the $k$-index matricization of a tensor $\mathbf{A}\in\mathbb{R}^{K_{1}\times\dots\times K_{N}}$ (see, e.g., hackbusch2012tensor for a definition) and let
$\widehat{q}_{i}=\mbox{rank}(\mbox{unfold}_{i}(\boldsymbol{\Phi})),\quad\text{and}\quad\widehat{r}_{k}=\mbox{rank}(\mbox{matr}_{k}(\boldsymbol{\Phi})),$ (30)
for $i=1,\dots,\text{\small$C+D$}$, $k=1,\dots,\text{\small$C+D$}-1$, and $\widehat{r}_{0}=\widehat{r}_{C+D}=1$. In other words, $\{\widehat{q}_{1},\dots,\widehat{q}_{C+D}\}$ and $\{\widehat{r}_{0},\dots,\widehat{r}_{C+D}\}$ are the HOSVD and TT-SVD ranks of $\boldsymbol{\Phi}$, respectively.
For the purpose of analysis, it is convenient to think about TT decomposition as a tensor chain. For example, for the decomposition (26) we define $r_{k-1}^{\mathbf{i}}\times N_{k}\times r_{k}^{\mathbf{i}}$ tensors $\mathbf{G}_{\mathbf{i}}^{(k)}$ by $\mathbf{G}_{\mathbf{i}}^{(k)}=\mathbf{g}^{\mathbf{i}}_{k,:,:}$ and consider the representation of $\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}$ via the chain of tensors,
$\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}\sim\left\{\mathbf{G}_{\mathbf{i}}^{(1)},\dots,\mathbf{G}_{\mathbf{i}}^{(D)}\right\}.$ (31)
Proposition 1
Tensor $\boldsymbol{\Phi}$ can be represented in the HTT format with $C$-ranks $\{q_{1},\dots,q_{C}\}$ and $D$-ranks $\{r_{0}^{(\mathbf{i})},\dots,r_{D}^{(\mathbf{i})}\}$, $\mathbf{i}\in\Omega_{C}^{q}$}, so that it holds
$q_{i}=\widehat{q}_{i},\quad\text{and}\quad r_{0}^{(\mathbf{i})}=1,~r_{j}^{(\mathbf{i})}=\widehat{r}_{j+C},$ (32)
for $i=1,\dots,\text{\small$C$}$, $j=1,\dots,\text{\small$D$}$, and for all $\mathbf{i}\in\Omega_{C}^{q}$.
Proof
Let $\{\mathbf{C},\mathrm{U}^{(1)},\dots,\mathrm{U}^{(C+D)}\}$ be the HOSVD decomposition of $\boldsymbol{\Phi}$ with the core tensor $\mathbf{C}$. We note the identity
$\mbox{matr}_{k}(\boldsymbol{\Phi})=\left(\mathrm{U}^{(k)}\otimes\dots\otimes\mathrm{U}^{(1)}\right)\mbox{matr}_{k}(\mathbf{C})\left(\mathrm{U}^{(C+D)}\otimes\dots\otimes\mathrm{U}^{(k+1)}\right)^{T},$ (33)
where the Kronecker tensor products of $\mathrm{U}^{(i)}$ are the matrices with orthonormal columns. This implies $\mbox{rank}(\mbox{matr}_{k}(\boldsymbol{\Phi}))=\mbox{rank}(\mbox{matr}_{k}(\mathbf{C}))$ and hence the TT-SVD ranks of $\mathbf{C}$ and $\boldsymbol{\Phi}$ are the same. Further, let
$\mathbf{C}\sim\left\{\mathbf{G}^{(1)},\cdots,\mathbf{G}^{(C+D)}\right\}$ (34)
be the TT-SVD decomposition of $\mathbf{C}$, where $\mathbf{G}^{(j)}$ are order three tensors of sizes $\widehat{r}_{j-1}\times\widehat{q}_{j}\times\widehat{r}_{j}$, and define $\widehat{\mathbf{G}}^{(j)}=\mathbf{G}^{(j)}\times_{2}\mathrm{U}^{(j)}$ for $j=\text{\small$C$}+1,\dots,\text{\small$C+D$}$. Then, an HTT decomposition of $\boldsymbol{\Phi}$ is given by the matrices $\{\mathrm{U}^{(1)},\dots,\mathrm{U}^{(C)}\}$ and the TT tensors
$\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}\sim\left\{\widetilde{\mathbf{G}}^{(C+1)}_{\mathbf{i}},\widehat{\mathbf{G}}^{(C+2)},\cdots,\widehat{\mathbf{G}}^{(C+D)}\right\},$ (35)
where $\widetilde{\mathbf{G}}^{(C+1)}_{\mathbf{i}}=\widehat{\mathbf{G}}^{(C+1)}\times_{1}\mathbf{v}_{\mathbf{i}}$ and $\mathbf{v}_{\mathbf{i}}\in\mathbb{R}^{\widehat{r}_{C}}$ is a vector defined through the product of matrices: $\mathbf{v}_{\mathbf{i}}=\mathbf{G}^{(1)}_{:,i_{1},:}\cdots\mathbf{G}^{(C)}_{:,i_{C},:}$.
The HTT representation guaranteed by Proposition 1 does not necessarily have minimal HTT ranks. In particular, in the representation (35), only $\widetilde{\mathbf{G}}^{(C+1)}$ depends on $\mathbf{i}$, and for all $\mathbf{i}$, the constructed tensors $\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}$ have the TT representation with the same ranks, which can be non-optimal. Moreover, we note that the HTT decomposition, by its construction, is independent of the position of the sliced indices in the tensor $\boldsymbol{\Phi}$. We consider the first $C$ for convenience. However, the TT ranks of $\boldsymbol{\Phi}$ do depend on the permutation of indices. Therefore, the result of Proposition 1 can be improved to achieve the minimum of TT ranks over all such permutations of modes in $\boldsymbol{\Phi}$ that preserve the order of the $D$-modes.
2.5 Accuracy of HTT completion
In this section, we prove an estimate for the accuracy of HTT completion in terms of the accuracy of the three steps for computing $\widetilde{\boldsymbol{\Phi}}$: sampling, projection, and component-wise completion.
We recall that $\mathrm{U}^{(i)}$, $i=1,\dots,C$, are orthogonal matrices of left singular vectors of $i$-th mode unfoldings of the tensor $\boldsymbol{\Phi}^{P}$, cf. (14) and (15). Denote the orthogonal projection on the column space of $\mathrm{U}^{(i)}$ by $\mathrm{P}_{(i)}=\mathrm{U}^{(i)}\big(\mathrm{U}^{(i)}\big)^{T}$ and define
$\varepsilon_{s}=\max\limits_{i=1,\ldots,C}\frac{\|\mathrm{P}_{(i)}^{\perp}\Phi^{(i)}\|_{F}}{\|\Phi^{(i)}\|_{F}},\quad\text{with}~\Phi^{(i)}=\mbox{unfold}_{i}(\boldsymbol{\Phi}),$ (36)
where $\mathrm{P}_{(i)}^{\perp}=\mathrm{I}-\mathrm{P}_{(i)}$. The values of $\varepsilon_{s}$ quantify how representative is the sampling of $\boldsymbol{\Phi}$, with smaller values corresponding to better representation.
Next, the accuracy of the projected tensor $\boldsymbol{\Phi}^{q}$ depends on the threshold $\varepsilon_{C}\geq 0$ used in (17) to determine the $q$-ranks. Finally, we assume that during the component-wise completion the tensors $\boldsymbol{\Phi}_{\mathbf{i}}^{q}:=\boldsymbol{\Phi}_{\mathbf{i},:}^{q}$ are reconstructed with the accuracy $\varepsilon_{q}\geq 0$, i.e., the bound
$\|\boldsymbol{\Phi}_{\mathbf{i}}^{q}-\widetilde{\boldsymbol{\Phi}}_{\mathbf{i}}^{q}\|_{F}\leq\varepsilon_{q}\|\boldsymbol{\Phi}_{\mathbf{i}}^{q}\|_{F}$ (37)
holds uniformly for all $\mathbf{i}\in\Omega_{C}^{q}$.
The following lemma provides the estimate of the completion accuracy.
Proposition 2
The following estimate holds
$\|\boldsymbol{\Phi}-\widetilde{\boldsymbol{\Phi}}\|_{F}\leq\big(\sqrt{C}(\varepsilon_{C}+\varepsilon_{s})+\varepsilon_{q})\big)\|\boldsymbol{\Phi}\|_{F},$ (38)
with $\varepsilon_{C}$, $\varepsilon_{s}$, and $\varepsilon_{q}$ from (17), (36), and (37), respectively.
Proof
By the triangle inequality we have
$\|\boldsymbol{\Phi}-\widetilde{\boldsymbol{\Phi}}\|_{F}\leq\|\boldsymbol{\Phi}-\boldsymbol{\Phi}^{q}\times_{1}\widetilde{\mathrm{U}}^{(1)}\ldots\times_{C}\widetilde{\mathrm{U}}^{(C)}\|_{F}+\|(\boldsymbol{\Phi}^{q}-\widetilde{\boldsymbol{\Phi}}^{q})\times_{1}\widetilde{\mathrm{U}}^{(1)}\ldots\times_{C}\widetilde{\mathrm{U}}^{(C)}\|_{F}.$ (39)
Using the definition of the Frobenious norm, its invariance under the tensor-matrix products with orthogonal $\widetilde{\mathrm{U}}^{(i)}$’s, and (37), we estimate the second term on the right hand side:
$\|(\boldsymbol{\Phi}^{q}-\widetilde{\boldsymbol{\Phi}}^{q})\times_{1}\widetilde{\mathrm{U}}^{(1)}\ldots\times_{C}\widetilde{\mathrm{U}}^{(C)}\|_{F}^{2}$ $=\|\boldsymbol{\Phi}^{q}-\widetilde{\boldsymbol{\Phi}}^{q}\|_{F}^{2}=\sum_{\mathbf{i}\in\Omega_{C}^{q}}\|\boldsymbol{\Phi}^{q}_{\mathbf{i}}-\widetilde{\boldsymbol{\Phi}}^{q}_{\mathbf{i}}\|_{F}^{2}$ (40)
$\leq\varepsilon_{q}^{2}\sum_{\mathbf{i}\in\Omega_{C}^{q}}\|\boldsymbol{\Phi}_{\mathbf{i}}^{q}\|_{F}^{2}=\varepsilon_{q}^{2}\|\boldsymbol{\Phi}\|_{F}^{2}.$
To handle the first term on the right hand side of (39), define auxiliary tensor $\widehat{\boldsymbol{\Phi}}=\boldsymbol{\Phi}\times_{1}{\mathrm{P}}_{(1)}\ldots\times_{C}\mathrm{P}_{(C)}$. It is easy to see that $\boldsymbol{\Phi}^{q}\times_{1}\widetilde{\mathrm{U}}^{(1)}\ldots\times_{C}\widetilde{\mathrm{U}}^{(C)}$ is the truncated HOSVD of $\widehat{\boldsymbol{\Phi}}$ and so it satisfies (see (de2000multilinear, , Property 10)) the bound
$\|\widehat{\boldsymbol{\Phi}}-\boldsymbol{\Phi}^{q}\times_{1}\widetilde{\mathrm{U}}^{(1)}\ldots\times_{C}\widetilde{\mathrm{U}}^{(C)}\|_{F}\leq\sqrt{C}\varepsilon_{C}\|\widehat{\boldsymbol{\Phi}}\|_{F}.$ (41)
Since $\widehat{\boldsymbol{\Phi}}$ is an orthogonal projection of $\boldsymbol{\Phi}$ along the first $C$ modes, it also holds $\|\widehat{\boldsymbol{\Phi}}\|_{F}\leq\|\boldsymbol{\Phi}\|_{F}$. Hence by the triangle inequality we obtain
$\|\boldsymbol{\Phi}-\boldsymbol{\Phi}^{q}\times_{1}\widetilde{\mathrm{U}}^{(1)}\ldots\times_{C}\widetilde{\mathrm{U}}^{(C)}\|_{F}\leq\|\boldsymbol{\Phi}-\widehat{\boldsymbol{\Phi}}\|_{F}+\sqrt{C}\varepsilon_{C}\|\boldsymbol{\Phi}\|_{F}.$ (42)
Consider now the decomposition
$\boldsymbol{\Phi}-\widehat{\boldsymbol{\Phi}}=\boldsymbol{\Phi}\times_{1}{\mathrm{P}}_{(1)}^{\perp}+(\boldsymbol{\Phi}\times_{2}{\mathrm{P}}_{(2)}^{\perp})\times_{1}{\mathrm{P}}_{(1)}+(\boldsymbol{\Phi}\times_{3}{\mathrm{P}}_{(3)}^{\perp})\times_{1}{\mathrm{P}}_{(1)}\times_{2}{\mathrm{P}}_{(2)}+\ldots$ (43)
$+(\boldsymbol{\Phi}\times_{C}{\mathrm{P}}_{(C)}^{\perp})\times_{1}{\mathrm{P}}_{(1)}\ldots\times_{C-1}\mathrm{P}_{(C-1)}.$
The terms in this decomposition are mutually orthogonal, as can be seen from the identities
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\langle\mathbf{A}\times_{i}{\mathrm{P}}_{(i)}^{\perp},\mathbf{B}\times_{i}{\mathrm{P}}_{(i)}\rangle_{F}}$ $=\langle{\mathrm{P}}_{(i)}^{\perp}A_{(i)},{\mathrm{P}}_{(i)}B_{(i)}\rangle_{\ell^{2}}$ (44)
$=\mbox{tr}\left({\mathrm{P}}_{(i)}^{\perp}A_{(i)}B_{(i)}^{T}{\mathrm{P}}_{(i)}\right)=\mbox{tr}\left({\mathrm{P}}_{(i)}{\mathrm{P}}_{(i)}^{\perp}A_{(i)}B_{(i)}^{T}\right)=0,$
which hold for two tensors $\mathbf{A}$ and $\mathbf{B}$ of the same sizes as $\boldsymbol{\Phi}$ and their unfoldings along the $i$th mode $A_{(i)}$ and $B_{(i)}$, respectively. Employing this decomposition, the orthogonality property and (36), we obtain the bound
$\|\boldsymbol{\Phi}-\widehat{\boldsymbol{\Phi}}\|_{F}^{2}$ $=\|\boldsymbol{\Phi}\times_{1}{\mathrm{P}}_{(1)}^{\perp}\|_{F}^{2}+\|(\boldsymbol{\Phi}\times_{2}{\mathrm{P}}_{(2)}^{\perp})\times_{1}{\mathrm{P}}_{(1)}\|_{F}^{2}+\ldots$ (45)
$\qquad\qquad+\|(\boldsymbol{\Phi}\times_{C}{\mathrm{P}}_{(C)}^{\perp})\times_{1}{\mathrm{P}}_{(1)}\ldots\times_{C-1}\mathrm{P}_{(C-1)}\|_{F}^{2}$
$\leq\|\boldsymbol{\Phi}\times_{1}{\mathrm{P}}_{(1)}^{\perp}\|_{F}^{2}+\|\boldsymbol{\Phi}\times_{2}{\mathrm{P}}_{(2)}^{\perp}\|_{F}^{2}+\ldots+\|\boldsymbol{\Phi}\times_{C}{\mathrm{P}}_{(C)}^{\perp}\|_{F}^{2}$
$=\|\mathrm{P}_{(1)}^{\perp}\Phi^{(1)}\|_{F}^{2}+\|\mathrm{P}_{(2)}^{\perp}\Phi^{(2)}\|_{F}^{2}+\ldots+\|\mathrm{P}_{(C)}^{\perp}\Phi^{(C)}\|_{F}^{2}$
$\leq\varepsilon_{s}^{2}\sum_{i=1}^{C}\|\Phi^{(i)}\|_{F}^{2}=C\varepsilon_{s}^{2}\|\boldsymbol{\Phi}\|_{F}^{2},$
where we used the fact that the Frobenious norm of a tensor equals to the Frobenious norm of any unfolding of it. Finally, the bound in (38) follows from (39), (40), (42), and (45).
2.6 Adaptive completion.
In many applications, and in particular for building ROM for parametric dynamical systems with accessible error bounds, one is interested in controlling the completion error, i.e., in ensuring that the estimate
$\|\boldsymbol{\Phi}-\widetilde{\boldsymbol{\Phi}}\|_{F}\leq\varepsilon\|\boldsymbol{\Phi}\|_{F},$ (46)
holds for a desired $\varepsilon\geq 0$. Such a bound is provided by the result in (38). In Algorithm 1 we can set $\varepsilon_{C}$ and have a control of $\varepsilon_{q}$ depending on the fitting algorithm used for pointwise tensor completion. However $\varepsilon_{s}$ is implicitly determined by the the index set $\widetilde{\Omega}_{D}$. This motivates an adaptive slice sampling completion algorithm, which ensures that a proxy of (46) is fulfilled. For the adaptive completion we define a testing set of indexes ${\Omega}_{D}^{\rm test}\subset\Omega_{D}$ and gradually increase the training set $\widetilde{\Omega}_{D}$ up to the point when the completed tensor satisfies
$\|\boldsymbol{\Phi}|_{\Omega^{\rm test}}-\widetilde{\boldsymbol{\Phi}}|_{\Omega^{\rm test}}\|_{F}\leq\varepsilon\|\boldsymbol{\Phi}|_{\Omega^{\rm test}}\|_{F},\quad\text{with}~\Omega^{\rm test}=\Omega_{C}\otimes{\Omega}_{D}^{\rm test}.$ (47)
Assuming that for any given $\mathbf{j}\in\Omega_{D}$ we may retrieve a slice $\boldsymbol{\Phi}_{:,\mathbf{j}}$ of $\boldsymbol{\Phi}$, the adaptive completion algorithm can be summarized as follows.
Algorithm 2 (Adaptive slice sampling tensor completion)
Input: Testing set ${\Omega}_{D}^{\rm test}\subset\Omega_{D}$, initial training set $\widetilde{\Omega}_{D}\subset\Omega_{D}$, threshold $\varepsilon$, maximum number of steps $N_{\max}$, and increments $P_{i}\in\mathbb{N}_{+}$, $i=1,\ldots,N_{\max}$.
  1. 1.
    For $\mathbf{j}\in\widetilde{\Omega}_{D}$ retrieve slices $\boldsymbol{\Phi}_{:,\mathbf{j}}$ to form the training data $\mathcal{D}$.
  2. 2.
    For $\mathbf{j}\in{\Omega}_{D}^{\rm test}$ retrieve slices $\boldsymbol{\Phi}_{:,\mathbf{j}}$ to compute $\boldsymbol{\Phi}|_{\Omega^{\rm test}}$.
  3. 3.
    Let $\varepsilon_{C}=\varepsilon/\sqrt{C}$.
  4. 4.
    For $i=1,\ldots,N_{\max}$:
    • (a)
      Execute Algorithm 1 with inputs $\widetilde{\Omega}_{D}$, $\mathcal{D}$, and $\varepsilon_{C}$;
    • (b)
      Break if (47) holds;
    • (c)
      Add $P_{i}$ more indices to $\widetilde{\Omega}_{D}$, update $\mathcal{D}$.
[…]
3 ROM for parametric dynamical systems
3.1 Projection TROM
For fast and accurate computations of trajectories $\mathbf{u}(t,\boldsymbol{\alpha})$ solving (1) for any $\boldsymbol{\alpha}\in\mathcal{A}$, we consider a projection based ROM which uses the proposed HTT LRTD instead of the conventional POD as the dimension reduction technique. The HTT decomposition allows for the recovery of information about the parameter dependence of reduced spaces from a smaller set of pre-computed snapshots. This information is exploited for building parameter specific ROMs. The LRTD–ROM was introduced in mamonov2022interpolatory and further developed and analyzed in mamonov2024tensorial ; mamonov2023analysis . However, in those earlier works it is assumed that a full snapshot tensor $\boldsymbol{\Phi}$ is available and therefore its LRTD in one of the standard tensor rank revealing formats (e.g., CP, Tuker, and TT) can be computed.
For the sake of exposition, we assume that all spatial degrees of freedom are vectorized into a single mode, so that only space and time remain as physical modes; i.e., $C=2$ for the rest of this section. In projection TROM we aim to find a low dimensional subspace $V^{\ell}(\boldsymbol{\alpha})\in\mathbb{R}^{M_{1}}$ with $\ell=\dim(V^{\ell}(\boldsymbol{\alpha}))$, which is parameter-specific and an approximation $\mathbf{u}^{\rm rom}(t,\boldsymbol{\alpha})$ to $\mathbf{u}(t,\boldsymbol{\alpha})$ is found by solving (1) projected onto $V^{\ell}(\boldsymbol{\alpha})$. This TROM solution $\mathbf{u}^{\rm rom}(t,\boldsymbol{\alpha})$ is then given by its coordinates in an orthogonal basis for $V^{\ell}(\boldsymbol{\alpha})$.
In turn, the orthogonal basis of $V^{\ell}(\boldsymbol{\alpha})$ is recovered by its coordinates in the basis of the universal reduced space $\widetilde{U}$, which is the span of all first-mode fibers of the low-rank tensor:
$\widetilde{U}=\text{range}\big(\widetilde{\Phi}_{(1)}\big).$ (48)
By the construction of $\widetilde{\boldsymbol{\Phi}}$, an orthogonal basis for $\widetilde{U}$ is given by the columns of the matrix $\mathrm{U}^{(1)}$ from (16).
In an interpolatory version of the TROM (see olshanskii2024approximating for a non-interpolatory version), one finds $V^{\ell}(\boldsymbol{\alpha})$ through interpolation procedure for the slices of $\widetilde{\boldsymbol{\Phi}}$. To define it, we first consider an interpolation operator that approximates a smooth function $g:\mathcal{A}\to\mathbb{R}$ using its values at the grid nodes $^{\widehat{\mathcal{A}}}$. More precisely, we assume ${\boldsymbol{\chi}}^{i}\,:\,\mathcal{A}\to\mathbb{R}^{K_{i}},\quad i=1,\dots,D,$ such that for any continuous function $g:\mathcal{A}\to\mathbb{R}$,
$I(g):=\sum_{k_{1}=1}^{K_{1}}\dots\sum_{k_{D}=1}^{K_{D}}\left({\boldsymbol{\chi}}^{1}(\boldsymbol{\alpha})\right)_{k_{1}}\dots\left({\boldsymbol{\chi}}^{D}(\boldsymbol{\alpha})\right)_{k_{D}}g\big(\widehat{\alpha}_{1}^{k_{1}},\dots,\widehat{\alpha}_{D}^{k_{D}}\big)$ (49)
defines an interpolant for $g$. One straightforward choice is the Lagrange interpolation of order $p$: for any $\boldsymbol{\alpha}\in\mathcal{A}$, let $\widehat{\alpha}_{i}^{i_{1}},\ldots,\widehat{\alpha}_{i}^{i_{p}}$ be the $p$ closest grid nodes to $\alpha_{i}$ on $\alpha_{i}^{\min},\alpha_{i}^{\max}]$, for $i=1,\ldots,D$. Then,
$\big({\boldsymbol{\chi}}^{i}(\boldsymbol{\alpha})\big)_{j}=\begin{cases}\prod\limits_{\begin{subarray}{c}m=1,\\ m\neq k\end{subarray}}^{p}(\widehat{\alpha}_{i}^{i_{m}}-\alpha_{i})\Big/\prod\limits_{\begin{subarray}{c}m=1,\\ m\neq k\end{subarray}}^{p}(\widehat{\alpha}_{i}^{i_{m}}-\widehat{\alpha}_{i}^{j}),&\text{if }j=i_{k}\in\{i_{1},\ldots,i_{p}\},\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad 0,&\text{otherwise},\end{cases}$ (50)
are the entries of ${\boldsymbol{\chi}}^{i}(\boldsymbol{\alpha})$ for $j=1,\dots,K_{i}$.
Given ${\boldsymbol{\chi}}^{i}$, we introduce the ‘local’ low-rank matrix $\widetilde{\Phi}(\boldsymbol{\alpha})$ via the in-tensor interpolation procedure for tensor $\widetilde{\boldsymbol{\Phi}}$:
$\widetilde{\Phi}(\boldsymbol{\alpha})=\widetilde{\boldsymbol{\Phi}}\times_{3}{\boldsymbol{\chi}}^{1}(\boldsymbol{\alpha})\times_{4}{\boldsymbol{\chi}}^{2}(\boldsymbol{\alpha})\dots\times_{D+2}{\boldsymbol{\chi}}^{D}(\boldsymbol{\alpha})\in\mathbb{R}^{M_{1}\times M_{2}}.$ (51)
[…]
$V^{\ell}(\boldsymbol{\alpha})=\mbox{range}\left([\mathrm{S}(\boldsymbol{\alpha})]_{1:\ell}\right),$ (52)
where
$\widetilde{\Phi}(\boldsymbol{\alpha})=\mathrm{S}(\boldsymbol{\alpha})\Sigma(\boldsymbol{\alpha})\mathrm{V}(\boldsymbol{\alpha})^{T},$ (53)
is the SVD of $\widetilde{\Phi}(\boldsymbol{\alpha})$.
3.2 In-tensor interpolation and finding the local basis
A remarkable fact is that we do not need to build the matrix $\widetilde{\Phi}(\boldsymbol{\alpha})$ and compute its SVD to find an orthogonal basis for $V^{\ell}(\boldsymbol{\alpha})$ once $\widetilde{\boldsymbol{\Phi}}$ is given in the HTT format. To see this, assume $\widetilde{\boldsymbol{\Phi}}$ is given in the HTT format with $C$-ranks $q_{1}$ and $q_{2}$ and define the parameter-dependent core matrix $\mathrm{C}_{\chi}(\boldsymbol{\alpha})\in\mathbb{R}^{q_{1}\times q_{2}}$ as follows
$\mathrm{C}_{\chi}(\boldsymbol{\alpha})]_{i_{1},i_{2}}=\widetilde{\boldsymbol{\Phi}}_{\mathbf{i}}^{q}\times_{1}{\boldsymbol{\chi}}^{1}(\boldsymbol{\alpha})\dots\times_{D}{\boldsymbol{\chi}}^{D}(\boldsymbol{\alpha}),\quad\text{with}~\mathbf{i}=(i_{1},i_{2}).$ (54)
Practically, the tensor matrix products in (54) are computed as follows. Each $\mathbf{g}^{\mathbf{i}}_{j}$ from (26) is a $\widetilde{r}^{\mathbf{i}}_{j-1}\times K_{j}\times\widetilde{r}^{\mathbf{i}}_{j}$ tensor, then $G^{\mathbf{i}}_{j}=\mathbf{g}^{\mathbf{i}}_{j}\times_{2}{\boldsymbol{\chi}}^{j}(\boldsymbol{\alpha})$ is a $\widetilde{r}^{\mathbf{i}}_{j-1}\times\widetilde{r}^{\mathbf{i}}_{j}$ matrix and we calculate $\mathrm{C}_{\chi}(\boldsymbol{\alpha})]_{i_{1},i_{2}}=G^{\mathbf{i}}_{1}\dots G^{\mathbf{i}}_{D}$ as a product of $D$ small size matrices.
By the definition of $\widetilde{\Phi}(\boldsymbol{\alpha})$ and (23) we have
$\begin{split}\widetilde{\Phi}(\boldsymbol{\alpha})&=\widetilde{\boldsymbol{\Phi}}\times_{3}{\boldsymbol{\chi}}^{1}(\boldsymbol{\alpha})\dots\times_{D+2}{\boldsymbol{\chi}}^{D}(\boldsymbol{\alpha})\\ &=(\widetilde{\boldsymbol{\Phi}}^{q}\times_{1}\widetilde{\mathrm{U}}^{(1)}\times_{2}\widetilde{\mathrm{U}}^{(2)})\times_{3}{\boldsymbol{\chi}}^{1}(\boldsymbol{\alpha})\dots\times_{D+2}{\boldsymbol{\chi}}^{D}(\boldsymbol{\alpha})\\ &=(\widetilde{\boldsymbol{\Phi}}^{q}\times_{3}{\boldsymbol{\chi}}^{1}(\boldsymbol{\alpha})\dots\times_{D+2}{\boldsymbol{\chi}}^{D}(\boldsymbol{\alpha}))\times_{1}\widetilde{\mathrm{U}}^{(1)}\times_{2}\widetilde{\mathrm{U}}^{(2)}\\ &=\mathrm{C}_{\chi}(\boldsymbol{\alpha})\times_{1}\widetilde{\mathrm{U}}^{(1)}\times_{2}\widetilde{\mathrm{U}}^{(2)}=\widetilde{\mathrm{U}}^{(1)}\mathrm{C}_{\chi}(\boldsymbol{\alpha})(\widetilde{\mathrm{U}}^{(2)})^{T}.\end{split}$ (55)
We combine this representation of $\widetilde{\Phi}(\boldsymbol{\alpha})$ with the SVD of the core matrix
$\mathrm{C}_{\chi}(\boldsymbol{\alpha})=\mathrm{U}_{c}\Sigma_{c}\mathrm{V}_{c}^{T}$ (56)
to obtain
$\widetilde{\Phi}(\boldsymbol{\alpha})=\left(\widetilde{\mathrm{U}}^{(1)}\mathrm{U}_{c}\right)\Sigma_{c}\left(\widetilde{\mathrm{U}}^{(2)}\mathrm{V}_{c}\right)^{T}.$ (57)
The right-hand side of (57) is the thin SVD of $\widetilde{\Phi}(\boldsymbol{\alpha})$, since the matrices $\widetilde{\mathrm{U}}^{(1)}$, $\mathrm{U}_{c}$, $\widetilde{\mathrm{U}}^{(2)}$, and $\mathrm{V}_{c}$ are all orthogonal. We conclude that the coordinates $\left\{\boldsymbol{\beta}_{1}(\boldsymbol{\alpha}),\dots,\boldsymbol{\beta}_{\ell}(\boldsymbol{\alpha})\right\}$ of the local reduced basis in the orthogonal basis of the universal space $\widetilde{U}$ are the first $\ell$ columns of $\mathrm{U}_{c}$. The parameter-specific ROM basis is then $\{\mathbf{z}_{i}(\boldsymbol{\alpha})\}_{i=1}^{\ell}$, with $\mathbf{z}_{i}(\boldsymbol{\alpha})=\widetilde{\mathrm{U}}^{(1)}\boldsymbol{\beta}_{i}(\boldsymbol{\alpha})$.
Note that the explicit computation of the basis functions $\mathbf{z}_{i}(\boldsymbol{\alpha})$ is not required during the online stage. In particular, the projection of the dynamical system onto the space $V^{\ell}(\boldsymbol{\alpha})=\mathrm{span}\{\mathbf{z}_{i}(\boldsymbol{\alpha})\}_{i=1}^{\ell}$ is performed using only the matrix of basis coordinates $\widetilde{\mathrm{U}}^{(1)}$; see Algorithm 3. Therefore, the essential information about $\widetilde{\boldsymbol{\Phi}}$ needed for the online part includes only the set of TT-tensors $\widetilde{\boldsymbol{\Phi}}_{\mathbf{i}}$:
$\mbox{online}(\widetilde{\boldsymbol{\Phi}})=\left\{\widetilde{\boldsymbol{\Phi}}_{\mathbf{i}},~~\mathbf{i}\in\Omega^{q}_{C}\right\}.$ (58)
3.3 Summary of the completion based TROM
Here we summarize the two stage completion based TROM (CTROM) approach. The core of the offline stage is finding a low-rank approximation of the snapshot tensor in the HTT format through the adaptive slice sampling completion Algorithm 2. Computing one slice of $\boldsymbol{\Phi}$ amounts to integrating (1) numerically for a given fixed parameter $\widehat{\boldsymbol{\alpha}}\in{\widehat{\mathcal{A}}}$, a task performed by the numerical solver $\mathcal{S}:\boldsymbol{\alpha}\to\Phi(\boldsymbol{\alpha})$. The second step of the offline stage is to project the system (1) onto the universal space $\widetilde{U}$, as defined in (48). The orthogonal basis for $\widetilde{U}$ is stored offline, while the projected system is passed to online stage. The particular computational details of the projection process depend on both the system form and its numerical solver $\mathcal{S}$. Detailed examples of system projections can be found mamonov2022interpolatory for the two particular systems studied in Section 4.
At the online stage, for any parameter $\boldsymbol{\alpha}\in\mathcal{A}$ one computes the orthogonal basis of the local reduced space $V^{\ell}(\boldsymbol{\alpha})$ represented by its coordinates in $\widetilde{U}$ and the system is projected the second time onto $V^{\ell}(\boldsymbol{\alpha})$. The computations at the online stage operate only with objects of reduced dimensions.
We summarize the two stage CTROM approach outlined above in Algorithm 3.
Algorithm 3 (Completion based TROM)
  • Offline stage.
    Input:
    • Numerical solver for (1) that computes the snapshot matrix (2) for a given $\boldsymbol{\alpha}\in\mathcal{A}$, i.e., $\mathcal{S}:\boldsymbol{\alpha}\to\Phi(\boldsymbol{\alpha})$;
    • Parameter grid $^{\widehat{\mathcal{A}}}$ as in (4);
    • Target accuracy $\varepsilon>0$;
    • Testing set ${\Omega}_{D}^{\rm test}\subset\Omega_{D}$, initial training set $\widetilde{\Omega}_{D}\subset\Omega_{D}$, maximum number of steps $N_{\max}$, increments $P_{i}\in\mathbb{N}_{+}$, $i=1,\ldots,N_{\max}$.
    1. 1.
[…]
    2. 2.
[…]
  • Online stage.
[…]
    1. 1.
[…]
    2. 2.
      Compute the SVD of the core matrix $\mathrm{C}_{\chi}(\boldsymbol{\alpha})=\mathrm{U}_{c}\Sigma_{c}\mathrm{V}_{c}^{T}$ with $\mathrm{U}_{c}=[\widetilde{\mathbf{u}}_{1},\ldots,\widetilde{\mathbf{u}}_{q_{\min}}]$;
    3. 3.
      Set $\boldsymbol{\beta}_{i}(\boldsymbol{\alpha})=\widetilde{\mathbf{u}}_{i}$, $i=1,\ldots,\ell$;
    4. 4.
      Project the solver $\widetilde{\mathcal{S}}$ onto the local reduced basis with coordinates $\{\boldsymbol{\beta}_{i}(\boldsymbol{\alpha})\}_{i=1}^{\ell}$ to obtain the parameter-specific CTROM solver $\widetilde{S}_{\ell}(\boldsymbol{\alpha})$.
    Output: Coordinates of the reduced basis in $\widetilde{U}$: $\{\boldsymbol{\beta}_{i}(\boldsymbol{\alpha})\}_{i=1}^{\ell}\subset\mathbb{R}^{q_{1}}$ and the CTROM solver $\widetilde{S}_{\ell}(\boldsymbol{\alpha})$.
3.4 Complexity of CTROM
The computational requirements of the completion-based TROM are determined by the following steps:
  • (i)
    Offline stage: Completion of $\widetilde{\boldsymbol{\Phi}}$ in the HTT format;
  • (ii)
[…]
  • (iii)
    Online stage: Computation of the coordinates of the parameter-specific reduced basis for a given $\boldsymbol{\alpha}$;
  • (iv)
    Solution of (1) projected onto the reduced space.
Since step (iv) is common to all projection-based ROMs, we focus on steps (i)–(iii). The offline stage is dominated by the completion algorithm and repeated FOM solves used to generate the training set. The number of FOM calls scales at least as $O(\sum_{j=1}^{D}r^{\rm max}_{j}K_{j}r^{\rm max}_{j+1})$, where $r^{\rm max}_{j}=\max_{\mathbf{i}\in\Omega_{C}^{q}}r^{\mathbf{i}}_{j}$. The cost of the completion step depends on the chosen scheme; in the numerical experiments we employ the stable rank-adaptive ALS (SALSA) method from grasedyck2019stable , with complexity $O(\sum_{j=1}^{D}|\Omega_{C}^{q}||\widetilde{\Omega}_{D}||r^{\rm max}_{j}r^{\rm max}_{j+1}|^{2})$ per adaptive step.
[…]
In step (iii), the cost of computing the $\boldsymbol{\alpha}$-specific reduced basis is governed by the interpolation procedure and the evaluation of the leading $n$ left singular vectors of the core matrix. Since the vectors $\mathbf{e}^{i}(\boldsymbol{\alpha})$ are sparse (typically $p=2$ or $3$ nonzero entries in Cartesian sampling), forming the core matrix $\mathrm{C}(\boldsymbol{\alpha})$ requires $O\Big(\sum_{i=2}^{D}\widetilde{r}_{i-1}\widetilde{r}_{i}\widetilde{r}_{i+1}\Big)$ operations. The subsequent SVD involves only the small $\widetilde{r}_{1}\times\widetilde{r}_{D+1}$ matrix. If a reduced basis in the physical space is required, the vectors are recovered as linear combinations of the columns of $\mathrm{U}$, with cost $O(|\Omega_{C}^{q}|\widetilde{r}_{1}\ell)$.
4 Numerical examples
We assess the performance of the sliced sampled completion and the completion based tensor ROM numerically in two examples of parametric dynamical systems. To address the component-wise completion problems for $\boldsymbol{\Phi}^{q}_{\mathbf{i}}$, we use SALSA method in the TT format from grasedyck2019stable , which ensures (37) with prescribed $\varepsilon_{q}$. All other input parameters of SALSA are left at default values. The corresponding compression $D$-ranks are determined adaptively during the completion process. We note that SALSA method does not necessarily produce a minimal-rank TT tensor that fits the data within the prescribed tolerance, and thus it can be viewed as an approximate solution to the completion problem (7).
The code of Matlab implementation of CTROM used for the numerical experiments is available upon request.
4.1 Parameterized heat equation
The first example is a dynamical system corresponding to the heat equation
$u_{t}(\mathbf{x},t,\boldsymbol{\alpha})=\Delta u(\mathbf{x},t,\boldsymbol{\alpha}),\quad\mathbf{x}\in\Omega,\quad t\in(0,T)$ (59)
in a rectangular domain with three holes $\Omega=\Omega_{r}\setminus(\Omega_{1}\cup\Omega_{2}\cup\Omega_{3})$ with $\Omega_{r}=[0,10]\times[0,4]$, as shown in Figure 1. Zero initial condition is enforced and the terminal time is set to $T=20$. The system has $D=4$ parameters that enter the boundary conditions:
$\left.(\mathbf{n}\cdot\nabla u+\alpha_{1}(u-1)\,)\right|_{\Gamma_{o}}$ $=$ $0,$ (60)
$\left.\left(\mathbf{n}\cdot\nabla u+\frac{1}{2}u\right)\right|_{\partial\Omega_{j}}$ $=$ $\frac{1}{2}\alpha_{j+1},\quad j=1,2,3,$ (61)
$\left.(\mathbf{n}\cdot\nabla u)\right|_{\partial\Omega_{r}\setminus\Gamma_{o}}$ $=$ $0,$ (62)
where, $\mathbf{n}$ is the outer unit normal and $\Gamma_{o}=0\times[0,4]$ is the left side of the rectangle $\Omega_{r}$. The parameter domain is the box $\mathcal{A}=[0.01,0.501]\times[0,0.9]^{3}$. The system (59)–(62) is discretized with $P_{2}$ finite elements on a quasi-uniform triangulation of $\Omega$ with maximum element size $h$.
Figure 1: Domain $\Omega$ and the solution $u(\mathbf{x},T,\boldsymbol{\alpha})$ of (59)–(62) corresponding to $\boldsymbol{\alpha}=(0.5,0,0,0.9)^{T}$.
In the numerical experiments below we assess the performance of CTROM for the system (59)–(62). For this purpose we employ the discrete $L^{2}(0,T;L^{2}(\Omega))$ norm error for a prescribed parameter value $\boldsymbol{\alpha}$, denoted by
$E_{\boldsymbol{\alpha}}=\Delta t\sum_{n=1}^{N}\left\|u(\;\cdot\;,t_{n},\boldsymbol{\alpha})-u_{\ell}(\;\cdot\;,t_{n},\boldsymbol{\alpha})\right\|^{2}_{L^{2}(\Omega)},$ (63)
where $u$ is the true (numerical) solution of (59)–(62), while $u_{\ell}$ is the CTROM solution. In order to evaluate CTROM accuracy over the whole parameter domain $\mathcal{A}$ we choose a testing set ${\mathcal{B}}=\{\widetilde{\boldsymbol{\alpha}}_{k}\}_{k=1}^{\widetilde{K}}\subset\mathcal{A}$ and compute the following maximum and mean quantities
${E}_{\max}$ $=$ $\left(\max_{k\in\{1,2,\ldots,\widetilde{K}\}}E_{\widetilde{\boldsymbol{\alpha}}_{k}}\right)^{1/2},$ (64)
${E}_{\text{mean}}$ $=$ $\left(\frac{1}{\widetilde{K}}\sum_{k=1}^{\widetilde{K}}E_{\widetilde{\boldsymbol{\alpha}}_{k}}\right)^{1/2}.$ (65)
To form $\mathcal{B}$} we choose $\widetilde{\boldsymbol{\alpha}}_{k}\in\mathcal{A}$ at random.
Experiment #\# 11 22 33 44 55
$h$ $0.33$ $0.2$ $0.1$ $0.2$ $0.2$
$K_{1},K_{2},K_{3},K_{4}] [$8,5,5,5][8,5,5,5] $[8,5,5,5]$ $[8,5,5,5]$ $[10,7,7,7]$ $[12,9,9,9]$
$|\widetilde{\Omega}_{D}|/|\Omega_{D}|$ $0.22$ $0.22$ $0.22$ $0.094$ $0.048$
$q_{1},q_{2}] [$51,18][51,18] $[51,18]$ $[51,18]$ $[50,18]$ $[51,18]$
$r_{1}^{\max},r_{2}^{\max},r_{3}^{\max}] [$7,8,5][7,8,5] $[7,7,5]$ $[7,8,5]$ $[8,7,7]$ $[7,7,7]$
$r_{1}^{\text{mean}}$ $6.002$ $6.003$ $6.025$ $6.008$ $6.005$
$r_{2}^{\text{mean}}$ $5.012$ $5.022$ $5.085$ $5.020$ $5.015$
$r_{3}^{\text{mean}}$ $4.014$ $4.026$ $4.090$ $4.046$ $4.042$
$\ell$ $18$ $18$ $18$ $18$ $18$
${E}_{\max}$ $2.729\cdot 10^{-3}$ $2.764\cdot 10^{-3}$ $2.779\cdot 10^{-3}$ $1.256\cdot 10^{-3}$ $4.937\cdot 10^{-4}$
${E}_{\text{mean}}$ $29\cdot 10^{-3} 1.331$$31\cdot 10^{-3}$ $1.338\cdot 10^{-3}$ $4.512\cdot 10^{-4}$ $2.484\cdot 10^{-4}$
Table 1: Results of numerical experiments for the heat equation.
We perform five numerical experiments, numbered 151-5 with parameters and results reported in Table 1. The following quantities were used in all experiments: $N=100$ time-domain snapshots were computed; threshold value $\varepsilon=\sqrt{2}\cdot 10^{-6}$ was set in Algorithm 2 so that the corresponding threshold in Algorithm 1 takes the value of $\varepsilon_{C}=10^{-6}$. The rank-adaptive componenent-wise TT completion for $\boldsymbol{\Phi}^{q}_{i}$, $\mathbf{i}\in\Omega_{C}^{q}$ was ran to ensure (37) with $\varepsilon_{q}=10^{-6}$.
For each experiment we report the $C$-ranks $q_{1},q_{2}]$ from step 2(c) of Algorithm 1 as well as the following quantities from the the last iteration of Algorithm 2: the sampling rate $|\widetilde{\Omega}_{D}|/|\Omega_{D}|$, the maximum and mean $D$-ranks
$r_{k}^{\max}=\max\limits_{\mathbf{i}\in\Omega_{C}^{q}}r_{k}^{\mathbf{i}},\quad r_{k}^{\text{mean}}=\frac{1}{|\Omega_{C}^{q}|}\sum_{\mathbf{i}\in\Omega_{C}^{q}}r_{k}^{\mathbf{i}},\quad k=1,2,\ldots,D-1,$ (66)
from TT completions in step 3(b) of Algorithm 1.
In experiments 1-3 the sampling set $\widehat{\mathcal{A}}$ was fixed with $K_{1}=8$, $K_{2}=5$, $K_{3}=5$ and $K_{4}=5$ grid nodes in each parameter direction, respectively. At the same time, the spatial FEM discretization mesh was refined with the maximum element size decreasing from $h=0.33$ to $h=0.2$ and then to $h=0.1$. We observe in the corresponding columns of Table 1 that CTROM is robust with respect to the mesh refinement in all of the reported quantities. This includes the $C$-ranks $q_{1},q_{2}]$, the maximum and mean $D$-ranks (66), as well as both error quantities (64)–(65) that barely change during mesh refinement.
In experiments 2, 4, 5 the FEM mesh is kept the same with $h=0.2$, but the set $\widehat{\mathcal{A}}$ is refined from $8\times 5^{3}=1000$ to $10\times 7^{3}=3430$ and then to $12\times 9^{3}=8748$ grid nodes, respectively. The purpose of this set of experiments is to show that the sampling rate required for achieving the same level of accuracy decreases as the size of $\widehat{\mathcal{A}}$ grows. Indeed, the sampling rate $|\widetilde{\Omega}_{D}|/|\Omega_{D}|$ first decreases from $0.22$ to $0.094$ and then to $0.048$. At the same time, the $C$-ranks $q_{1},q_{2}]$ and the maximum and mean $D$-ranks (66) remain stable, while the error quantities (64)–(65) improve (decrease). This demonstrates that slice sampling tensor completion alleviates the curse of dimensionality in that the sampling rate can be decreased for larger $\widehat{\mathcal{A}}$ without the loss of accuracy of the resulting CTROM.
In the final experiment for the parameterized heat problem, we investigate
Figure 2: CTROM solution error versus completion accuracy.
the dependence of the ROM solution error on the completion accuracy. The ROM solution error is defined in equations (64)–(65). The completion error is characterized by $\varepsilon$ in (46), and the estimate (38) provides an upper bound for it. Since we do not have direct access to $\varepsilon$ or to certain quantities appearing in (38), the completion accuracy is instead assessed on a test set, as described in (47). The results presented in Fig. 2 show that the ROM errors decrease almost linearly over the given range of $\varepsilon$. In this experiment, we set $h=0.2$, used a $10\times 7^{3}$ mesh in $\widehat{\mathcal{A}}$, and chose $\ell=q_{2}$.
4.2 Parameterized advection-diffusion equation
The second numerical example involves more (up to $D_{\max}=12$) parameters than that in Section 4.1 and corresponds to the dynamical system resulting from the discretization of a linear advection-diffusion equation
$u_{t}(\mathbf{x},t,\boldsymbol{\alpha})=\nu\Delta u(\mathbf{x},t,\boldsymbol{\alpha})-\boldsymbol{\eta}(\mathbf{x},\boldsymbol{\alpha})\cdot\nabla u(\mathbf{x},t,\boldsymbol{\alpha})+f(\mathbf{x}),\quad\mathbf{x}\in\Omega,\quad t\in(0,T)$ (67)
in the unit square domain $\Omega=[0,1]\times[0,1]\subset\mathbb{R}^{2}$, $\mathbf{x}=(x_{1},x_{2})^{T}\in\Omega$, with final time $T=1$. Here $\nu=1/30$ is the diffusion coefficient, $\boldsymbol{\eta}:\Omega\times\mathcal{A}\to\mathbb{R}^{2}$ is the parameterized advection field and $f(\mathbf{x})$ is a Gaussian source
$f(\mathbf{x})=\frac{1}{2\pi\sigma_{s}^{2}}\exp\left(-\cfrac{(x_{1}-x_{1}^{s})^{2}+(x_{2}-x_{2}^{s})^{2}}{2\sigma_{s}^{2}}\right),$ (68)
with $\sigma_{s}=0.05$, $x_{1}^{s}={x_{2}^{s}}=0.25$. Homogeneous Neumann boundary conditions and zero initial condition are imposed
$\left.\left(\mathbf{n}\cdot\nabla u\right)\right|_{\partial\Omega}=0,\quad u(\mathbf{x},0,\boldsymbol{\alpha})=0.$ (69)
The model parameters enter the divergence free advection field $\boldsymbol{\eta}$ defined as
$\boldsymbol{\eta}(\mathbf{x},\boldsymbol{\alpha})=\begin{pmatrix}\eta_{1}(\mathbf{x},\boldsymbol{\alpha})\\ \eta_{2}(\mathbf{x},\boldsymbol{\alpha})\end{pmatrix}=\begin{pmatrix}\cos(\alpha_{1})\\ \sin(\alpha_{1})\end{pmatrix}+\frac{1}{\pi}\begin{pmatrix}\partial_{x_{2}}h(\mathbf{x},\boldsymbol{\alpha})\\ -\partial_{x_{1}}h(\mathbf{x},\boldsymbol{\alpha})\end{pmatrix}$, (70)
where $h(\mathbf{x})$ is the cosine trigonometric polynomial
$\begin{split}h(\mathbf{x},\boldsymbol{\alpha})=&\;\;\;\;\alpha_{2}\cos(\pi x_{1})+\alpha_{3}\cos(\pi x_{2})+\alpha_{4}\cos(\pi x_{1})\cos(\pi x_{2})\\ &+\alpha_{5}\cos(2\pi x_{1})+\alpha_{6}\cos(2\pi x_{2})\\ &+\alpha_{7}\cos(2\pi x_{1})\cos(\pi x_{2})+\alpha_{8}\cos(\pi x_{1})\cos(2\pi x_{2})\\ &+\alpha_{9}\cos(2\pi x_{1})\cos(2\pi x_{2})+\alpha_{10}\cos(3\pi x_{1})+\alpha_{11}\cos(3\pi x_{2})\\ &+\alpha_{12}\cos(3\pi x_{1})\cos(\pi x_{2}).\end{split}$ (71)
The system (67)–(69) is discretized similarly to (59)–(62), but using a uniform grid in $\Omega$.
We perform the numerical experiments for the advection-diffusion system for varying number of parameters $D$ by setting $\alpha_{j}=0$ for $j=D+1,\ldots,12$. Then, the vectors $\alpha_{1},\ldots,\alpha_{D}]^{T}$ belong to the parameter domain that is the box $\mathcal{A}=[0.1\pi,0.3\pi]\times[-0.1,0.1]^{D-1}$.
$D$ 66 99 $12$
$h$ $\frac{1}{32}$ $\frac{1}{32}$ $\frac{1}{32}$
$K_{1}$ $10$ $10$ $10$
$K_{j}$, $j=2,\ldots,D$ 55 55 55
$|\widetilde{\Omega}_{D}|/|\Omega_{D}|$ $0.05056$ $0.00204$ $6\times 10^{-5}$
$q_{1},q_{2}] [$55,11][55,11] $[66,11]$ $[83,11]$
$\ell$ $11$ $11$ $11$
${E}_{\max}$ $1.890\cdot 10^{-4}$ $2.022\cdot 10^{-4}$
${E}_{\text{mean}}$ $1.242\cdot 10^{-4}$ $1.334\cdot 10^{-4}$
Table 2: Results of numerical experiments for the advection-diffusion equation with $D=6,9,12$.
We set $\varepsilon=5\cdot 10^{-4}$ in Algorithm 2 and $\varepsilon_{C}=10^{-6}$ in Algorithm 1. The rank-adaptive componenent-wise TT completion for $\boldsymbol{\Phi}^{q}_{i}$, $\mathbf{i}\in\Omega_{C}^{q}$ was performed to ensure (37) with $\varepsilon_{q}=10^{-4}$. The results of numerical experiments with parametrized advection–diffusion equation are summarized in Tables 23 and Figures 35.
Experiment #\# 11 22 33 44 55 66
$h$ $\frac{1}{32}$ $\frac{1}{32}$ $\frac{1}{32}$ $\frac{1}{32}$ $\frac{1}{64}$ $\frac{1}{16}$
$K_{1}$ $10$ 1919 $29$ $39$ $10$ $10$
$K_{j}$, $j=2,\ldots,6$ 55 99 $14$ 1919 55 55
$|\widetilde{\Omega}_{D}|/|\Omega_{D}|$ $0.05056$ $0.00262$ $0.00038$ $0.00008$ $0.05056$ $0.05056$
$q_{1},q_{2}] [$55,11][55,11] $[54,11]$ $[53,11]$ $[53,11]$ $[55,11]$ $[56,11]$
$r_{1}^{\text{mean}}$ $8.500$ $7.173$ $7.759$ $7.593$ $8.540$ $8.553$
$r_{2}^{\text{mean}}$ $12.800$ $10.227$ $11.855$ $11.236$ $12.727$ $12.823$
$r_{3}^{\text{mean}}$ $15.014$ $12.144$ $13.065$ $13.264$ $15.011$ $14.995$
$r_{4}^{\text{mean}}$ $15.805$ $18.075$ $19.807$ $20.001$ $15.809$ $15.797$
$r_{5}^{\text{mean}}$ $5.000$ $9.000$ $12.380$ $13.121$ $5.000$ $5.000$
$\ell$ $11$ $11$ $11$ $11$ $11$ $11$
${E}_{\max}$ $1.890\cdot$ $2.364\cdot$ $1.644\cdot$ $4\cdot 2.2772.2$$7\cdot$ $1.653\cdot$
$\cdot 10^{-4}$ $\cdot 10^{-4}$ $\cdot 10^{-4}$ $\cdot 10^{-4}$ $\cdot 10^{-4}$ $\cdot 10^{-4}$
${E}_{\text{mean}}$ $1.242\cdot$ $1.025\cdot$ $9.080\cdot$ $9.533\cdot$ $1.493\cdot$ $1.093\cdot$
$\cdot 10^{-4}$ $\cdot 10^{-4}$ $\cdot 10^{-5}$ $\cdot 10^{-5}$ $\cdot 10^{-4}$ $\cdot 10^{-4}$
Table 3: Results of numerical experiments for the advection-diffusion equation with $D=6$.
In the first series of experiments, we vary the parameter space dimension $D$, while keeping the discretization parameters fixed: $h=1/32$, $\Delta t=\frac{1}{200}$. The grid in the parameter domain is the Cartesian product of uniform grids for each parameter. We use 10 uniformly distributed nodes for $\alpha_{1}$ and 5 nodes for $\alpha_{i}$, $i>1$. The recovered $C$-ranks are given in Table 2 and the mean $D$-ranks were found to be [1, 8.50, 12.8, 15.0,15.8, 5.0, 1] for $D=6$, [1, 8.27, 14.2, 18.0, 20.8, 27.0, 29.7, 20.1, 5.0, 1] for $D=9$ and [1, 7.54, 12.6, 16.7, 20.8, 26.9, 30.6, 38.6, 48.9, 50.9, 25.0, 5.0, 1] for $D=12$. We see that both ranks increase with $D$, reflecting the growing variability of the solution for more complex advection fields as more parameters are involved. For $D=12$, however, we were only able to achieve a completion accuracy of $\epsilon=10^{-3}$, which required fitting approximately $27\cdot 10^{3}$ parameters from the training set. The adaptive completion algorithm we used was not efficient for a larger number of training parameters in this example.
Figure 3: Completion statistics versus completion accuracy for parameter space dimensions $D=6,9,12$. Left panel: Percent of the observed tensor elements; Central panel: Compression factor; Right panel: Total number of degrees of freedom for representation in HTT format.
Figure 4: Completion statistics versus completion accuracy for various spatial resolutions. Left panel: Percent of the observed tensor elements; Central panel: Compression factor; Right panel: Total number of degrees of freedom for representation in HTT format.
Figure 5: Completion statistics versus completion accuracy for various resolutions of the parameter domain. Left panel: Percent of the observed tensor elements; Central panel: Compression factor as a ratio in HTT format; Right panel: A rescaled total number of degrees of freedom for representation in HTT format.
The percentage of observed entries of the snapshot tensor $\boldsymbol{\Phi}$ required for its completion depends on $D$ and the targeted accuracy $\varepsilon$; see the left panel in Figure 3. It decreases from $5\%$ for $D=6$ to $0.006\%$ for $D=12$ as required for completion with $\varepsilon=5\times 10^{-4}$ accuracy; see the ratio $|\widetilde{\Omega}_{D}|/|\Omega_{D}|$ in Table 2.
Other completion statistics presented in Figure 3 include the compression factor and the total degrees of freedom (d.o.f.) required for representation in the HTT format. The compression factor, defined as the ratio between the number of d.o.f. in HTT and full tensor format, and the total number of d.o.f., grows as $\varepsilon$ decreases, as expected, reaching around $10^{6}$ times compression for $D=12$. For the total d.o.f. in HTT format, the observed rate of growth with $\varepsilon\to 0$ lies between $O(\varepsilon^{-\frac{1}{3}})$ and $O(\varepsilon^{-\frac{2}{3}})$; see Figure 3, right panel.
We now study how the completion statistics change when the discretization parameters vary. Specifically, we vary the space discretization parameter $h$ and refine the mesh in the parameter domain, keeping the number of parameters fixed ($D=6$). Table 3 shows that both $C$ and $D$ ranks are not sensitive to the refinement or coarsening of meshes in both physical and parameter domains.
[…]
The final experiment for the advection-diffusion equation is to compare the performance of CTROM to the conventional TROM that uses the full tensor of snapshots, as introduced in mamonov2022interpolatory . For this comparison the setup is as follows. We use $D=9$ parameters and the grid $^{\widehat{\mathcal{A}}}$ with $K_{1}=5$ and $K_{j}=3$, $j=2,\ldots,D$, for a total of $32,805$ grid points. While $^{\widehat{\mathcal{A}}}$ is rather coarse, it is still the finest grid that can be handled by the conventional TROM for $D=9$ on the machine used for the numerical experiments. We set $\varepsilon=5\cdot 10^{-4}$ in Algorithm 2 and we use the same accuracy for TT decomposition in the offline stage of the conventional TROM. We compare the two TROM aproaches by timing the offline stage of both, since it dominates the computational cost. For the conventional TROM the computation of the snapshots for the tensor $\boldsymbol{\Phi}$ took 2828hrs 5151min with additional 1414min to perform its TT decomposition for a total of 2929hrs 0505min for the offline stage. The offline stage of CTROM, Algorithm 2, took 77hrs 1010min which includes computing $4,172$ samples, i.e., the percentage of observed entries of $\boldsymbol{\Phi}$ is $12.71\%$. Thus, for this particular case we observe more than four-fold gain in computational cost of CTROM compared to the conventional TROM. Note also that the relatively high observed entry percentage in Algorithm 2 is due to a coarse grid $^{\widehat{\mathcal{A}}}$. One should expect even greater performance gains of CTROM for finer grids $^{\widehat{\mathcal{A}}}$.
We conclude this section by noting that the accuracy of the HTT-ROM is found to be insensitive to variations in dimensions and meshes (see the error statistics in Tables 2 and 3). It is determined by the accuracy of the tensor completion and $\ell$. We consistently use $\ell=11$, which is sufficiently large for the HTT-ROM error to be dominated by the completion error.
5 Conclusions
Motivated by the problem of reconstructing the parametric solution manifold of a dynamical system through sparse sampling, this paper introduces a low-rank tensor format for the inexact completion of a tensor observed slice-wise. The completion procedure is based on solving a sequence of “standard” tensor completion problems using a common format, with the tensor train format selected in this work. These standard completion problems can be addressed in parallel.
Applying the completion method to two linear parametric parabolic PDEs, discretized via the finite element method, revealed the following properties: (i) The resulting tensor ranks are insensitive to discretization parameters in space and time, as well as to grid resolution in the parameter domain. (ii) The ranks increase with more accurate completion, requiring additional memory for storing the recovered tensor. This dependency is of the form $\text{d.o.f.}=O(\varepsilon^{-\alpha})$, where $\varepsilon$ represents the completion accuracy and $\alpha\geq 0$ is an exponent. In our tests, $\alpha$ did not exceed $\frac{2}{3}$. (iii) The total number and percentage of observed tensor entries required for successful completion depends on the target accuracy, but more significantly on the dimension of the parameter space and on how fine is the parameter domain grid. Specifically, the total number of entries required increases roughly linearly with the number of nodes for any parameter. Therefore, if the grid in $\mathcal{A}$ is refined by doubling the nodes in each direction, the number of observed entries required doubles, while the percentage of observed entries ‘decrease’ by a factor of $2^{D-1}$. This explains our ability to recover snapshot tensors from fewer than $0.01\%$ of entries in higher-dimensional parameter spaces. (iv) The compression achieved by the HTT format for snapshot tensors similarly depends on parameter space dimension and resolution. In our numerical examples, we achieved a compression factor of $10^{6}$ while preserving an accuracy of $10^{-4}$.
In the context of parametric dynamical systems, we found the available rank-adaptive tensor completion method effective and efficient for problems with up to a dozen parameters. Extending this approach significantly beyond this number of parameters warrants further investigation.
Acknowledgements.
A.M. and M.O. were supported by the U.S. National Science Foundation under award DMS-2309197. This material is based upon research supported in part by the U.S. Office of Naval Research under award number N00014-21-1-2370 to A.M.